OpenStudy (anonymous):

In 1940 the average size of a U.S. farm was 174 acres. Let's say that the standard deviation was 55 acres. Suppose we randomly survey 38 farmers from 1940. Give the distribution of XBar xbar ~ N ( , )

OpenStudy (anonymous):

\[\bar{X}=\frac{1}{n}\sum_{i=1}^nX_i\] \(\bar{X}\) will be normally distributed with mean \(E(\bar{X})\) and variance \(V(\bar{X})\). Suppose \(E(X_i)=\mu\) and \(V(X_i)=\sigma^2\), then \[\begin{align*}E(\bar{X})&=E\left(\frac{1}{n}\sum_{i=1}^nX_i\right)\\ &=\frac{1}{n}E\left(\sum_{i=1}^nX_i\right)\\ &=\frac{1}{n}\sum_{i=1}^nE\left(X_i\right)\\ &=\frac{1}{n}\sum_{i=1}^n\mu\\ &=\frac{1}{n}(n\mu)\\\\ &=\mu\end{align*}\] \[\begin{align*}V(\bar{X})&=V\left(\frac{1}{n}\sum_{i=1}^nX_i\right)\\ &=\frac{1}{n^2}\sum_{i=1}^nV\left(X_i\right)\\ &=\frac{1}{n^2}(n\sigma^2)\\ &=\frac{\sigma^2}{n}\\ SD(\bar{X})&=\sqrt{\frac{\sigma^2}{n}}\\ &=\frac{\sigma}{\sqrt n} \end{align*}\] You're given that \(\mu=174\), \(\sigma=55\) and \(n=38\), so \[\bar{X}\sim N\left(\mu,~\frac{\sigma}{\sqrt n}\right)\]